OpenStudy (anonymous):

How do i find the Y intercept of 1/(0-3)^2 + 2

OpenStudy (anonymous):

There's no function. The constant you gave, 1/9+2, or 19/9.

OpenStudy (anonymous):

how did you get the answer though? i have a ton of this problems to do but i want to learn the right way to do it :( soooo confused

OpenStudy (anonymous):

You gave me a constant with no context. It's like me asking, "What's the y-intercept of 4?" There's no answer to that question.

OpenStudy (anonymous):

oh k well this was the full question. I only asked a part of it. Use transformations y= 1/x or y = 1/x^2 to sketch the rational function label all intercepts and asymptotes its asked to graph ( which i can do ) then it asks what is the Y - intercept f(x) = 1/(x-3)^2 + 2

OpenStudy (anonymous):

You made a typo in the original question; you put in 0 instead of x. Which, ironically, was what you are supposed to do. The y-intercept is at the y-axis, which is when x=0.

OpenStudy (anonymous):

i meant to do that ;) so for this equation would i do this... 3^2*2+1 = 19 3^2 = 9 so 19/9?

OpenStudy (anonymous):

f(x) = 1/(x-3)^2 + 2 was the equation, right? f(0) = 1/(0-3)^2 + 2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

f(0) = 1/(0-3)^2 + 2 this is your answer. 0 in for x